# DiverseNet: Decision Diversified Semi-supervised Semantic Segmentation Networks for Remote Sensing Imagery

### v1

@@ V-A Implementation Details
We implemented our methods by using the PyTorch framework. Following [chen2018encoder], we used a mini-batch SGD optimizer and adopted a poly learning rate policy, where the current learning rate equals the initial learning rate multiplied by $\left(1-\frac{\text{ iter }}{\text{ max-iter }}\right)^{\text{power }}$. The initial learning rate and power are set to 0.01 and 0.9, respectively. All experiments were conducted on the GW4 Isambard with an NVIDIA A100-sxm GPU and an AMD EPYC 7543P CPU  [gw4].
@@
We implemented various classic semi-supervised learning methods (MT, CCT, GCT and CPS) on remote sensing datasets. All these methods use DeepLabv3+ with ResNet-50 pre-trained on ImageNet as the semantic segmentation network. To be fair the proposed DiverseHead is also based on DeepLabv3+ with the pretrained ResNet-50.
@@
V-B Quantitative Results and Analysis
@@
We evaluated the performance of the proposed approach and reference networks using 4 aforementioned datasets, and the average results are presented in Table I. The reason we first share average results across datasets is to provide a global applicability display for all utilised techniques, and each data set’s performance will be discussed in the sequel. Evaluating the performance results in Table I, we conclude that the proposed DiverseModel achieves the best performance among all metrics. In particular, for metrics like UA and mIoU, the proposed DiverseModel exhibits an improvement of over 5% compared to another cross-model-based approach, CPS. Although the performance of DiverseHead is lower than that of DiverseModel, it outperforms other reference methods and reaches the second best. Upon convergence of contemporary methodologies with remote sensing datasets, MT demonstrates notably inferior average performance across various metrics, with the exception of UA. Both CCT and GCT exhibit comparable performance, yet the performance superiority of CCT is notably evident in its PA metric. CPS consistently yields superior outcomes across the majority of metrics when juxtaposed against other aforementioned state-of-the-art methodologies.
Meanwhile, it is worth noting that the DiverseHead consists of a notably lighter architecture compared to all other models where Table II presents the required number of parameters during training for each semi-supervised learning architecture. It is evident that the architectures using multiple segmentation networks during training typically need higher memory demands. In contrast, the proposed DiverseHead is remarkably lightweight as it only employs a single model with multiple heads which consists of only 2 convolution layers, eliminating the need for multiple models during training. Thus, the parameter size of the DiverseHead (DT&DF) is only 16% bigger than that of the single network (Base in Table II), whereas the parameter size of other reference semi-supervised architectures is at least twice that of the single network. Despite this need for a lower number of parameters, the DiverseHead proposed method outperforms all state-of-the-art with at least 1% accuracy whilst obtaining similar performance to the companion method of DiverseModel.
@@
Table I: Average performance comparison with the state of the arts on 4 datasets. DT and DF present Dropout and dynamic freezing, respectively.
OA UA PA mIoU $F_{1}$
MT[tarvainen2017mean] 85.63% 73.14% 80.10% 63.93% 76.39%
CCT[ouali2020semi] 86.60% 72.17% 83.29% 64.07% 77.22%
GCT[ke2020guided] 86.53% 73.33% 80.50% 64.59% 76.72%
CPS[chen2021semi] 87.17% 72.89% 84.62% 65.02% 78.18%
DiverseHead (DT) 88.42% 77.80% 84.95% 69.70% 81.05%
DiverseHead (DF) 88.29% 76.98% 84.87% 68.85% 80.58%
DiverseModel 88.86% 78.24% 86.06% 70.53% 81.79%
@@
Table II: The required parameter size in each semi-supervised learning approach. Base means supervised segmentation network DeepLabv3+. DT and DF present Dropout and dynamic freezing, respectively.
CPS [chen2021semi] MT [tarvainen2017mean] CCT [ouali2020semi] GCT [ke2020guided]
@@
Base DiverseHead(DT) DiverseHead(DF) DiverseModel
Size (MB) 151.689 175.806 175.806 936.262
Specifically, we present the results of each semi-supervised learning method for the 4 utilised datasets in Table III. DiverseNet architectures reach a remarkable performance for each of the datasets. Specifically, the proposed DiverseModel achieves the best performance on both RoadNet and Massachusetts building datasets for all metrics, and the DiverseHead (DT) shows the second best for the most of metrics on these two datasets. For the other two datasets, Potsdam and DFC2020, The DiverseHead (DF) method shows competitive and better performance compared to the companion DiverseModel for most of the metrics.
@@
Table III: Performance comparison with the state of the arts on 4 datasets. DT and DF present Dropout and dynamic freezing, respectively.
OA UA PA mIoU $F_{1}$
Potsdam MT[tarvainen2017mean] 81.98% 73.66% 78.39% 63.07% 75.95%
CCT[ouali2020semi] 82.66% 74.64% 77.62% 64.16% 76.10%
GCT[ke2020guided] 83.99% 75.65% 80.81% 65.80% 78.14%
CPS[chen2021semi] 85.00% 75.76% 82.94% 66.69% 79.19%
DiverseHead (DT) 84.60% 77.42% 80.49% 67.18% 78.92%
DiverseHead (DF) 85.37% 78.43% 81.86% 68.49% 80.11%
DiverseModel 85.76% 76.75% 83.45% 67.85% 79.96%
DFC2020 MT[tarvainen2017mean] 78.64% 59.59% 73.57% 50.44% 65.85%
CCT[ouali2020semi] 79.71% 59.87% 76.40% 51.04% 67.13%
GCT[ke2020guided] 80.84% 61.47% 71.43% 52.17% 66.07%
CPS[chen2021semi] 81.50% 61.75% 79.47% 53.21% 69.50%
DiverseHead (DT) 82.11% 62.31% 80.38% 53.93% 70.20%
DiverseHead (DF) 82.19% 62.29% 80.73% 54.02% 70.32%
DiverseModel 81.87% 62.20% 80.69% 53.71% 70.25%
RoadNet MT[tarvainen2017mean] 93.00% 80.74% 85.54% 72.96% 83.07%
CCT[ouali2020semi] 93.90% 83.85% 87.26% 76.23% 85.52%
GCT[ke2020guided] 93.57% 79.26% 89.25% 73.51% 83.96%
CPS[chen2021semi] 94.36% 84.21% 88.85% 77.44% 86.47%
DiverseHead (DT) 96.28% 90.89% 91.95% 84.84% 91.42%
DiverseHead (DF) 95.36% 87.99% 90.31% 81.35% 89.13%
DiverseModel 96.79% 92.82% 92.57% 86.87% 92.70%
Massachusetts MT[tarvainen2017mean] 88.92% 78.56% 82.90% 69.25% 80.67%
CCT[ouali2020semi] 89.65% 79.91% 84.23% 70.95% 82.01%
GCT[ke2020guided] 87.73% 76.95% 80.51% 66.88% 78.69%
CPS[chen2021semi] 87.85% 69.85% 87.24% 62.76% 77.58%
DiverseHead (DT) 90.69% 80.56% 86.97% 72.86% 83.64%
DiverseHead (DF) 90.25% 79.22% 86.59% 71.53% 82.74%
DiverseModel 91.01% 81.19% 87.53% 73.67% 84.24%
@@
We also considered a downgraded version of the current DiverseHead approach, called single-head supervision (SHS). The segmentation network consists of only a single head. During each iteration, the current model’s predictions for unlabeled data are utilized to generate pseudo labels through an $\arg\max$ operation. These generated pseudo labels are then used to calculate the unsupervised loss. The baseline is the segmentation network only trained by labelled data. The performance of the baseline and SHS, along with the proposed DiverseHead, are presented in Table IV in order to provide experimental evidence of the novelty of utilisation of multiple heads in DiverseHead architecture. While the single-head supervision benefits from the pseudo labels and exhibits better performance compared to the baseline model especially improving the metric of PA, both versions of DiverseHead approaches demonstrate a further significant improvement by taking advantage of multiple heads. Specifically, the mIoU of DiverseHead (DF) is 5.7% higher than that of Baseline and approximately 4% higher than that of SHS for Potsdam. Also, the mIoU value of DiverseHead (DT) is about 10% higher than that of SHS.
@@
Table IV: performance comparison with single-head supervision (SHS) and baseline model which is only supervised by labelled data. DT and DF present Dropout and dynamic freezing, respectively.
OA UA PA mIoU $F_{1}$
Potsdam Baseline 81.64% 73.86% 76.05% 62.72% 74.94%
@@
DiverseHead (DT) 84.60% 77.42% 80.49% 67.18% 78.92%
DiverseHead (DF) 85.37% 78.43% 81.86% 68.49% 80.11%
RoadNet Baseline 93.56% 82.85% 86.53% 75.06% 84.65%
SHS 93.79% 81.19% 88.62% 74.87% 84.74%
DiverseHead (DT) 96.28% 90.89% 91.95% 84.84% 91.42%
DiverseHead (DF) 95.36% 87.99% 90.31% 81.35% 89.13%
Table V: Ablation study about the number of heads for DiverseHead with dynamic freezing
@@
Potsdam 8 85.70% 78.47% 82.09% 68.82% 80.24%
10 85.37% 78.43% 81.86% 68.49% 80.11%
12 85.51% 77.23% 83.03% 68.14% 80.02%
RoadNet 8 94.28% 80.20% 92.23% 75.59% 85.79%
10 95.36% 87.99% 90.31% 81.35% 89.13%
12 95.35% 85.57% 92.22% 80.52% 88.77%
Average 8 89.99% 79.33% 87.16% 72.21% 83.02%
10 90.36% 83.21% 86.08% 74.92% 84.62%
12 90.43% 81.40% 87.62% 74.33% 84.40%
An ablation study was conducted on the Potsdam and RoadNet datasets to explore the impact of the number of heads for the proposed DiverseHead method with dynamic freezing perturbation. The results in Table V indicate that using 10 heads yields the optimal choice across most metrics in terms of average performance. The performance difference between the versions with different numbers of heads is very small on Potsdam, but the version of 8 heads shows a performance drop. Considering the stability, We chose the 10-head version of DiverseHead (DF) for all experiments in this paper.
Table VI: Performance comparison of the DiverseModel with its constituent networks
OA UA PA mIoU $F_{1}$
@@
Since the DiverseModel technique uses three different segmentation networks, we also evaluated the performance of each individual member network within the DiverseModel in Table VI. The results indicate that the cross-supervision of all networks shows a significant improvement in the segmentation task performance compared to each separate network.
Figure 4: Visual results of each method on Potsdam Dataset. Values in sub-captions refer to the IoU scores for each method of this case.
Figure 5: Visual results of each method on RoadNet Dataset. Values in sub-captions refer to the IoU scores for each method of this case.
Figure 6: Visual results of each method on DFC2020 Dataset. Values in sub-captions refer to the IoU scores for each method of this case.
Figure 7: Visual results of each method on Massachusetts Dataset. Values in sub-captions refer to the IoU scores for each method of this case.
V-C Qualitative Results and Analysis
The visible results are presented in Figures from 4 to 7. For each dataset, we chose one case and depicted the RGB image, ground truth, and predictions of all the methods. Compared to the ground truth, the proposed three methods (right-top 3 sub-figures) achieve better similarity to the ground truth compared to the reference models. We also calculated the IoU scores for all predictions with the ground truth for this case, which are shown in the sub-caption of each prediction. The IoU scores also demonstrate that the performance of the proposed DiverseNet methods is more competitive than that of the other approaches.
VI Conclusion
In this paper, we proposed a lightweight semi-supervised learning based on a multi-head structure called DiverseHead, which equipped a dynamic freezing and dropout perturbation approaches. The multi-head idea benefits from bagging and shows competitive performance for the segmentation of remote sensing imagery. Also, another high-efficiency multiple network based semi-supervised learning method called DiverseModel is investigated for further analysis. DiverModel shows the strongest performance on various datasets compared to the state of the arts but requires more computational resources, whilst the performance of DiverseHead is slightly lower than DiverseModel but dramatically lighter than all reference models.

### v2

@@ 6.1 Implementation Details
We implemented our methods by using the PyTorch framework. Following (Chen et al., 2018), we used a mini-batch SGD optimizer and adopted a poly learning rate policy, where the current learning rate equals the initial learning rate multiplied by $\left(1-\frac{\text{ iter }}{\text{ max-iter }}\right)^{\text{power }}$. The initial learning rate and power are set to 0.01 and 0.9, respectively. All experiments were conducted on the GW4 Isambard with an NVIDIA A100-sxm GPU and an AMD EPYC 7543P CPU  (McIntosh-Smith, 2014).
@@
We implemented various classic semi-supervised learning methods (MT, CCT, GCT and CPS) on remote sensing datasets. All these methods use DeepLabv3+ with ResNet-50 backbone pre-trained on ImageNet as the semantic segmentation network. To be fair the proposed DiverseHead is also based on DeepLabv3+ with the pretrained ResNet-50 on ImageNet.
@@
6.2 Quantitative Results and Analysis
@@
We evaluated the performance of the proposed approaches and some existing semi-supervised frameworks using the five performance metrics above. Experiments are conducted on the four previously mentioned datasets, namely Potsdam, DFC2020, RoadNet, and Massachusetts Building. The average results are presented in Table 1. The reason we first share average results across datasets is to provide a global performance for all utilised techniques, and the detailed performance comparison for all methods on each dataset will be discussed later. From the average results in Table 1, the proposed DiverseHead (DF) achieves the best overall performance in which there are 4 of 5 metrics reach the best compared to other methods, which are marked as red. DiverseModel shows the second-best performance among all methods in Table 1. In particular, for metrics of UA, the proposed DiverseModel exhibits an improvement of over 3.6% compared to another cross-model-based approach, CPS. Although the performance of DiverseHead (DT) is marginally lower than that of DiverseHead (DF) and DiverseModel, the three methods exhibit similar performance and surpass other compared methods. With these used remote sensing datasets, MT demonstrates notably inferior average performance across various metrics. CCT, GCT, and CPS exhibit comparable performance, yet the performance superiority of CPS is notably evident in its PA metric.
@@
Table 1: Average performance comparison with the state-of-the-art methods on four datasets. DT and DF indicate dropout and dynamic freezing, respectively.
Models OA UA PA mIoU $F_{1}$
MT (Tarvainen and Valpola, 2017) 86.33% 74.90% 81.47% 65.97% 77.95%
CCT (Ouali et al., 2020) 87.10% 76.61% 82.29% 67.71% 79.19%
GCT (Ke et al., 2020) 87.43% 75.53% 82.04% 67.19% 78.63%
CPS (Chen et al., 2021) 88.03% 75.90% 85.17% 68.10% 80.13%
DiverseModel (Ma et al., 2023) 88.77% 79.54% 85.18% 70.92% 82.02%
DiverseHead (DT) 88.51% 78.40% 84.84% 70.07% 81.29%
DiverseHead (DF) 88.97% 78.99% 85.82% 71.14% 82.08%
It is important to highlight that the DiverseHead is a significantly lightweight semi-supervised learning architecture aimed at segmentation tasks within remote sensing imagery. Table 2 presents the required number of parameters during training for each semi-supervised learning architecture. Apart from DiverseHead, all other semi-supervised learning approaches typically involve parameters exceeding three hundred megabytes during training. Aside from DiverseModel, all other methods are based on the same segmentation network namely DeepLabv3+. DiverseModel is the largest architecture although it provides competitive performance compared to other cross-model-based methods namely MT, CCT, GCT, and CPS as shown in Table 1. It is evident that the architectures using multiple segmentation networks during training typically need higher memory demands. In contrast, the proposed DiverseHead is remarkably lightweight as it only employs a single model with multiple heads which consists of only a few small convolutional layers, eliminating the need for multiple models during training. Thus, the parameter size of the DiverseHead (DT&DF) is only 16% bigger than that of the single network (Base in Table 2), whereas the parameter size of other reference semi-supervised architectures is at least twice that of the single network. Although just a a few parameters are required, the DiverseHead proposed method outperforms this state-of-the-art with at least 1% accuracy and 3% mIoU whilst obtaining similar performance to the companion method of DiverseModel.
@@
Table 2: The required parameter size in each semi-supervised learning approach. Base means supervised segmentation network DeepLabv3+. DT and DF present dropout and dynamic freezing, respectively.
CPS (Chen et al., 2021) MT (Tarvainen and Valpola, 2017) CCT (Ouali et al., 2020) GCT (Ke et al., 2020)
@@
DiverseModel (Ma et al., 2023) DiverseHead(DT) DiverseHead(DF) Base
Size (MB) 936.262 175.806 175.806 151.689
Specifically, we present the results of each semi-supervised learning method for the 4 utilised datasets in Table 3. The performance of the DiverseNet family, namely DiverseHead and DiverseModel, is noticeably superior to that of the other listed semi-supervised learning methods across all evaluated datasets, as indicated by the results. In particular, for the Potsdam and Massachusetts Building datasets, the proposed DiverseHead attains the highest performance across 4 out of 5 segmentation metrics whilst DiverseModel emerges as the second-best method based on its results. However, in the case of the RoadNet and DFC2020 datasets, DiverseModel reaches the best for most performance metrics, securing the top position whilst DiverseHead closely follows as the second-best method, despite their similar performance. The performance of DiverseHead (DT) consistently falls between that of DiverseModel and DiverseHead (DF) for most used datasets apart from DFC2020. This dataset provides very low-resolution labels, distinguishing it from other datasets. This could be one of the reasons why DiverseHead (DT) outperforms DiverseHead (DF) in DFC2020.
@@
Table 3: Performance comparison with the state of the arts on four datasets. DT and DF present Dropout and dynamic freezing, respectively.
Model OA UA PA mIoU $F_{1}$
Potsdam MT(Tarvainen and Valpola, 2017) 81.98% 73.66% 78.39% 63.07% 75.95%
CCT(Ouali et al., 2020) 82.66% 74.64% 77.62% 64.16% 76.10%
GCT(Ke et al., 2020) 83.99% 75.65% 80.81% 65.80% 78.14%
CPS(Chen et al., 2021) 85.00% 75.76% 82.94% 66.69% 79.19%
DiverseModel(Ma et al., 2023) 85.76% 76.75% 83.45% 67.85% 79.96%
DiverseHead (DT) 84.66% 77.04% 80.78% 67.12% 78.87%
DiverseHead (DF) 85.98% 79.15% 82.87% 69.63% 80.97%
DFC2020 MT(Tarvainen and Valpola, 2017) 78.64% 59.59% 73.57% 50.44% 65.85%
CCT(Ouali et al., 2020) 79.71% 59.87% 76.40% 51.04% 67.13%
GCT(Ke et al., 2020) 80.84% 61.47% 71.43% 52.17% 66.07%
CPS(Chen et al., 2021) 81.49% 61.74% 79.46% 53.20% 69.49%
DiverseModel(Ma et al., 2023) 81.87% 62.20% 80.69% 53.71% 70.25%
DiverseHead (DT) 82.02% 62.13% 80.61% 53.81% 70.18%
DiverseHead (DF) 81.78% 61.81% 80.21% 53.46% 69.82%
RoadNet MT(Tarvainen and Valpola, 2017) 94.53% 86.77% 87.94% 78.81% 87.35%
CCT(Ouali et al., 2020) 95.45% 89.15% 89.91% 81.97% 89.53%
GCT(Ke et al., 2020) 95.20% 86.27% 90.93% 80.36% 88.54%
CPS(Chen et al., 2021) 95.66% 88.84% 90.95% 82.46% 89.88%
DiverseModel(Ma et al., 2023) 96.84% 93.12% 92.58% 87.11% 92.84%
DiverseHead (DT) 96.16% 91.70% 90.98% 84.70% 91.33%
DiverseHead (DF) 96.68% 91.99% 92.73% 86.33% 92.36%
Massachusetts MT(Tarvainen and Valpola, 2017) 90.16% 79.58% 85.97% 71.55% 82.65%
CCT(Ouali et al., 2020) 90.59% 82.77% 85.23% 73.67% 83.98%
GCT(Ke et al., 2020) 89.67% 78.75% 85.01% 70.42% 81.76%
CPS(Chen et al., 2021) 89.95% 77.26% 87.33% 70.07% 81.98%
DiverseModel(Ma et al., 2023) 90.62% 86.07% 84.01% 75.00% 85.03%
DiverseHead (DT) 91.21% 82.73% 86.98% 74.67% 84.80%
DiverseHead (DF) 91.43% 83.00% 87.45% 75.16% 85.17%
@@
In order to prove the efficiency of the multi-head framework, we considered a downgraded version of the current DiverseHead approach, called single-head supervision (SHS). The segmentation framework consists of only a single head. During each training iteration, the current model’s predictions for unlabeled data are used to generate pseudo labels through an $argmax$ operation. These generated pseudo labels are then used to calculate the unsupervised loss. Also, we provide the performance of the segmentation network namely DeepLabv3+, called Base in Table 4, only trained by labelled data (part of each dataset) without using unlabelled data. The performance of the Base and SHS, along with the proposed DiverseHead, are presented in Table 4 to provide experimental evidence of the improvement and efficiency of the use of multiple heads in the DiverseHead architecture. While the single-head supervision benefits from the pseudo labels and exhibits better performance compared to the Base model, especially improving the metric of PA, both versions of DiverseHead approaches demonstrate a further significant improvement by taking advantage of multiple heads. Specifically, the mIoU of DiverseHead (DF) is 6.91% higher than that of Base and 5.12% higher than that of SHS for Potsdam. While the performance of DiverseHead (DT) is slightly lower than that of DiverseHead (DF), its performance is still much better than that of SHS and Base. Similarly, DiverseHead demonstrates superior performance on the RoadNet dataset compared to SHS and Base, particularly evident in the mIoU metric, where it exhibits a significant improvement of 6.22% over Base.
@@
Table 4: Performance comparison with single-head supervision (SHS) and baseline model which is only supervised by labelled data. DT and DF present dropout and dynamic freezing, respectively.
Model OA UA PA mIoU $F_{1}$
Potsdam Base 81.64% 73.86% 76.05% 62.72% 74.94%
@@
DiverseHead (DT) 84.66% 77.04% 80.78% 67.12% 78.87%
DiverseHead (DF) 85.98% 79.15% 82.87% 69.63% 80.97%
RoadNet Base 94.86% 88.22% 88.31% 80.11% 88.26%
SHS 95.32% 87.08% 90.82% 80.95% 88.91%
DiverseHead (DT) 96.16% 91.70% 90.98% 84.70% 91.33%
DiverseHead (DF) 96.68% 91.99% 92.73% 86.33% 92.36%
Table 5: Ablation study on the number of heads for DiverseHead with dynamic freezing
@@
Potsdam                       8 85.92% 77.86% 83.86% 68.91% 80.75%
                      10 85.98% 79.15% 82.87% 69.63% 80.97%
                      12 85.70% 78.54% 82.55% 68.96% 80.50%
RoadNet                       8 96.51% 91.36% 92.55% 85.67% 91.95%
                      10 96.68% 91.99% 92.73% 86.33% 92.36%
                      12 96.66% 92.12% 92.58% 86.31% 92.35%
Average                       8 91.22% 84.61% 88.20% 77.29% 86.35%
                      10 91.33% 85.57% 87.80% 77.98% 86.66%
                      12 91.18% 85.33% 87.57% 77.64% 86.42%
An ablation study was conducted on the Potsdam and RoadNet datasets to explore the impact of the number of heads for the proposed DiverseHead method with dynamic freezing perturbation. The results in Table 5 indicate that using 10 heads yields the optimal choice across most metrics in terms of average performance though the performance difference between the versions with different numbers of heads is small on both datasets. Thus, we chose the 10-head version of DiverseHead for all experiments in this paper.
Table 6: Performance comparison of the DiverseModel with its constituent networks on Potsdam dataset
Model OA UA PA mIoU $F_{1}$
@@
Since the DiverseModel technique uses three different segmentation networks, we also evaluated the performance of each member network within the DiverseModel on the Potsdam dataset in Table 6. The component models UNet, SegNet, and PSPNet use unlabeled data through the way of SHS. DiverseModel demonstrates a significant improvement across all performance metrics in the segmentation task compared to each network. In particular, the metric of PA experiences a notable improvement of 4.57% compared to the best-performing individual component model. This phenomenon can be attributed to the enhancement of pseudo-label diversity through cross-model supervision, resulting in a significant improvement in PA (recall) during the test phase. The findings suggest that the cross-supervision of different networks has the potential to achieve superior performance compared to the best-performing individual component.
Potsdam DFC2020 RoadNet Massachusetts
Image [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Ground Truth [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
MT [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
IoU = 39.38% IoU = 70.08% IoU = 79.30% IoU = 73.37%
CCT [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
IoU = 51.19% IoU = 78.21% IoU = 86.24% IoU = 75.97%
GCT [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
IoU = 40.35% IoU = 76.66% IoU = 80.96 IoU = 74.20%
Potsdam DFC2020 RoadNet Massachusetts
CPS Refer to caption Refer to caption Refer to caption Refer to caption
IoU = 50.21% IoU = 79.39% IoU = 84.92% IoU = 74.16%
DiverseModel Refer to caption Refer to caption Refer to caption Refer to caption
IoU = 58.76% IoU = 80.28% IoU = 94.67% IoU = 76.84%
DiverseHead (Dropout) Refer to caption Refer to caption Refer to caption Refer to caption
IoU = 57.46% IoU = 80.12% IoU = 93.01% IoU = 78.93%
DiverseHead (Dynamic Freezing) Refer to caption Refer to caption Refer to caption Refer to caption
IoU = 58.25% IoU = 80.31% IoU = 90.35% IoU = 78.17%
Figure 7: Visual results of each method on four datasets. The IoU score of prediction for each method is shown in its sub-caption.
6.3 Qualitative Results and Analysis
The visible results are presented in Figures 7. For each dataset, we randomly selected one case of the RGB image, its ground truth, and predictions for all the methods. Also, the IoU scores are calculated for all predictions with the ground truth for each case, which are shown in the sub-caption of each prediction. Based on the IoU values of visual predictions, the highest score is obtained by either DiverseModel or DiverseHead (DT&DF). Especially for the RoadNet dataset, both the DiverseModel and DiverseHead families achieve an IoU of over 90% surpassing other methods by at least 3.5%. Visually, the segmentation maps of DiverseModel and DiverseHead display better continuity and show better similarity to the ground truth than other methods.
7 Conclusion
In this paper, we proposed a lightweight semi-supervised learning approach based on a multi-head structure called DiverseHead. Based on the multi-head structure we provide two perturbation methods, namely dynamic freezing and dropout. Taking inspiration from the theory of bagging, a voting mechanism is proposed to generate beneficial pseudo-labels in the training stage. This simple and lightweight semi-supervised learning framework shows competitive performance for the segmentation of remote sensing imagery. Furthermore, we conducted further analysis and additional evaluation on the previously proposed multi-network-based semi-supervised learning method known as DiverseModel. Based on the results obtained from the aforementioned four remote sensing datasets, DiverseHead and DiverseModel demonstrate comparable performance while significantly outperforming various widely used semi-supervised learning frameworks. It is important to highlight that the training structure of DiverseHead is significantly lighter than that of DiverseModel and other state-of-the-art methods, while still achieving competitive performance. This implies that DiverseHead is a better option for those who lack high-memory computation resources.